In January 2016​, United Airlines​ (UAL) had a market capitalization of $20.67 billion, debt of $12.04 billion, and cash of $5.07 billion. United Airlines had revenues of $38.88 billion. Southwest Airlines​ (LUV) had a market capitalization of $27.47​billion, debt of $3.34 billion, cash of $3.14 ​billion, and revenues of $19.71 billion.

a. Compare the market​ capitalization-to-revenue ratio​ (also called the​ price-to-sales ratio) for United Airlines and Southwest Airlines.

b. Compare the enterprise​ value-to-revenue ratio for United Airlines and Southwest Airlines.

c. Which of these comparisons is more​ meaningful? Explain.
Answer #1

1: Market cap to Revenue ratio

UAL = 20.67/38.88 = 0.53

Southwest = 27.47/19.71 = 1.39

UAL has a lower ratio and so it is more attractive investment since the company yis able to generate higher sales with a lower Market cap.

2: EV to Revenue Ratio

UAL = (Market cap + Debt – Cash)/ Sales

= (20.67+12.04-5.07)/38.88

=0.71

Southwest = (27.47+3.34-3.14)/19.71 = 1.4

The UAL has a lower EV to Sales ratio which means that it is more attractive or undervalued.

3: EV to Sales ratio is more meaningful since it takes into account the value of debt which is not considered in case of Market Cap to Sales ratio.